Prove that (Z(sqrt(2)),+,×) is an integral domain.



Subject: Rings and fields


1
Expert's answer
2021-05-25T07:12:54-0400

Solution:

 Consider "\mathbb{Z}[\sqrt{2}]=\{x+y \sqrt{2} \mid x, y \in
Z\}" of the real numbers R.

Since the real numbers R form a ring, the subset "\mathbb{Z}[\sqrt{2}]" and forms ring if they are closed under the addition and multiplication which is directly checked. It is also integral domain since R as a field is an integral domain.

Alternatively: a nontrivial solution of "(x+\sqrt{2} y)(a+\sqrt{2} b)=0" gives for"x b+y a \neq 0" the equation "\sqrt{2}=-\frac{x a+2 y b}{x b+y a}" , a contradiction to the irrationality of "\sqrt{2}" . If "x b+y a=0" it follows that "x a+2 y b=0" . Since the determinant of this linear system seen as an equation for x and y is "2 b^{2}-a^{2}" , the only solution is "x=y=0" if "2 b^{2}-a^{2} \neq 0" . If "2 b^{2}-a^{2}=0" , the irrationality of "\sqrt{2} \implies a=b=0" . Thus either "x+\sqrt{2} y\ or\ a+\sqrt{2} b" is zero.